if y varies inversely as x^2 and y=2 when x=4, what is the constant of variation?
Answer by stanbon(75887)

if y varies inversely as x^2 and y=2 when x=4, what is the constant of variation?
-------------
y = k/x^2
Need to find "k".
--------------
2 = k/4^2
k = 32
-------------
Cheers,
Stan H.